OpenStudy (anonymous):

find the lung capacity of a 40year old woman who is 165cm tall

OpenStudy (lgbasallote):

is there a formula for this??

OpenStudy (jlastino):

0.041h-0.018a-2.69 a= age h= height 0.041(165)-0.018(40)-2.69 3.355 liters

OpenStudy (anonymous):

Since "the lung capacity of a 40-year-old woman who is 160 centimeters tall", we know that and . Plug these values in to get

OpenStudy (lgbasallote):

@Jlastino may i ask where you got "0.041h-0.018a-2.69" ? is it a formula? i would love to learn this too just in case...

OpenStudy (anonymous):

so this means that the lung capacity is about 18.99 liters.

OpenStudy (jlastino):

@lgbasallote I searched a similar question through my profile :)) it gave this polynomial

OpenStudy (anonymous):

0.14(160)-0.018-(40)-2.69 =22.4-0.72-2.69= 18.99

OpenStudy (anonymous):

hope you gott it...riight

OpenStudy (anonymous):

r u there geaneete?

OpenStudy (lgbasallote):

hmmm wikipedia gave me this equation though... 0.0275*AgeInYears + 0.0189*HeightInCentimetres - 2.6139 so many variations @_@ hahahaha

OpenStudy (anonymous):

great =D

OpenStudy (anonymous):

variations are too diffrent lol

OpenStudy (lgbasallote):

@geannette maybe your teacher gave you a fixed equation? coz we are having different equations...

OpenStudy (anonymous):

whats your answer geanette?

OpenStudy (anonymous):

?